let W be an inner product space and let w1 and w2 be vectors in W. suppose ll w1 ll =5^1/2 and ll w2 ll= 4 and the angle between w1 and w2 is pi/3. compute

  • <w1,w1> and <w2,w2>
  • <w1, w2>
  • <w1+ 2w2, 3w1-w2>
1
Expert's answer
2021-01-13T18:22:46-0500

Answer:


Given w1 and w2 are two vectors in W and

ll w1 ll =51/2______(1)

 ll w2 ll =4_________(2)


Angle between w1 and w2="\frac {π} 3".

To compute "\coloneq"

  1. <w1,w1> and <w2,w2>
  2. <w1, w2>
  3. <w1+ 2w2, 3w1-w2>


We know that,

"If u1 and u2 are two vectors and "\theta" be the angle between them, then

"\cos" "\theta" ="⟨u1, u2 ⟩⟩ ⟩ \over ∥u1, u2 ∥" where "\theta" is angle between w1 and w2

"\over"

where ∥u1∥= "\sqrt {<u1, u1>}"


and ∥u2∥= "\sqrt {<u2, u2>}"


1.

(a) <w1,w1> =? ; given  ll w1 ll ="\sqrt 5"

using inner product property,

ll w1 ll = "\sqrt {<w1, w1>}"

ll w1 ll2 = <w1,w1> (square both sides)

<w1,w1> =  ll w1 ll2

= ("\sqrt 5")2

=(51/2)2 (multiply "1 \over 2" by 2)

=5


Therefore, <w1,w1> =5


(b)  <w2,w2>=? ; given  ll w2 ll =4

Using,

ll w2 ll= "\sqrt {<u2, u2>}"

squaring both sides we get,

  ll w2 ll2 = <w2,w2>

Hence, <w2,w2>=  ll w2 ll2

 <w2,w2>= (4)2 = 16

Therefore,   <w2,w2> = 16


(2)To find <w1, w2>

we know,

 "\cos" "\theta" = "⟨w1, w2 ⟩ \over ∥w1, w2 ∥" where "\theta" is angle between w1 and w

Given, ll w1 ll = "\sqrt 5"

llw2 ll = 4

and "\theta" = "\frac {π} 3"

  "\cos"  "\frac {π} 3"= "<w1, w2> \over 
(\sqrt 5) (4)"  


"1 \over 2" = "<w1, w2> \over 
(\sqrt 5) (4)"    (  "\cos"  "\frac {π} 3"=  "1 \over 2")


<w1, w2> =  "1 \over 2" ("\sqrt 5") (4) (divide 4 by 2 to eliminate the half)

=2 "\sqrt 5"


Therefore, <w1, w2> = 2 "\sqrt 5" ____(3)


(3)

To find <w1+ 2w2, 3w1-w2>.

We know, according to property of inner product.

<αa+βb, γc+δd> = <αa, γc+δd> + <βb, γc+δd>

where α, β, γ and δ are real constants and a,b,c,d are vectors.


Hence, <αa+βb, γc+δd> = α<a, γc+δd> + β<b, γc+δd>

= α(<a, γc> + <a, δd>) + β (<b, γc> + <b, δd>)

α(γ<a, c> + δ<a, d>) + β (γ<b, c> + δ<b, d>)

=αγ<a, c> + δα<a, d>) + βγ<b, c> + δβ<b, d>)

Here,

<w1+ 2w2, 3w1-w2> = <w1, 3w1-w2> +  <2w2, 3w1-w2>

=1<w1, 3w1> + 1<w1, -w2> +  2<w2, 3w1> + 2 <w2, -w2>

=1(3)<w1, w1> + 1(-1)<w1, w2> +  2(3)<w2, w1> + 2(-1) <w2, w2>


Hence,

<w1+ 2w2, 3w1-w2>=

=3<w1, w1> -1<w1, w2> +  6<w2, w1> - 2<w2, w2>


From previous solutions;

<w1, w1> = ll w1 ll2

<w1, w2> = <w2, w1>


<w1+ 2w2, 3w1-w2>

=3 ll w1 ll2 - <w1, w2> +  6<w2, w1> - 2 ll w2 ll2

 =3 ll w1 ll2 - <w1, w2> +  6< w1,w2> - 2 ll w2 ll2

=3 ll w1 ll2 +  5< w1,w2> - 2 ll w2 ll2


From (1) , (2) , (3) we have

 ll w1 ll = "\sqrt 5"  , ll w2 ll = 4, and < w1,w2> = 2 "\sqrt 5" 


Therefore,

<w1+ 2w2, 3w1-w2> = 3("\sqrt 5")2 + 5(2"\sqrt 5")- 2(4)2

  = 3(5) + 10"\sqrt 5" - 2 (16)

= 15+ 10"\sqrt 5" - 32

=-17+10"\sqrt 5"


Hence, <w1+ 2w2, 3w1-w2> = -17+10"\sqrt 5"